# On braided simple extensions and braided non-semisimple near-group categories

### v1

@@ 4.2. Non-semisimple braided near-group categories
Theorem 1. Let $\mathcal{C}$ be a non-semisimple braided near-group category with generalized fusion rule $(G,r)$, then $r=0$. In particular, a non-semisimple braided near-group category is weakly integral.
@@
Let $\mathcal{C}$ be a braided non-semisimple near-group category with a generalized fusion rule $(G,r)$, and assume that $r>0$. By 4.2.2, we get that $\mathcal{D}$ is a symmetric subcategory of $\mathcal{C}$ and $Q\in\mathcal{D}^{\prime}$. By [D], since $\mathcal{D}$ is not semisimple, it is braid equivalent to $\textrm{sRep}(G\ltimes W,u)$, a category of finite supermodules of the supergroup $G\ltimes W$, with $u\in G$ a central element of order $2$ acting as the parity. Since $\mathcal{D}$ is pointed, the simple objects of $\mathcal{D}$ generate a fusion subcategory, which is braid equivalent to $\text{Vec}_{G\times\mathbb{Z}/2\mathbb{Z}}$ (the central element $u$ generates $\mathbb{Z}/2\mathbb{Z}$). In particular, there exist a simple object $S\in\mathcal{D}_{f}\subset\mathcal{D}$ such that $S\otimes S=\mathbbm{1}$ and $c_{S,S}=-\tau_{S,S}$ for $\tau_{X,Y}$ the flip morphism $\tau_{X,Y}:X\otimes Y\to Y\otimes X$.
The objects $S$ and $Q$ are simple objects, and $S$ is invertible, thus we have an isomorphism $\psi:Q\otimes S\simeq Q$, and the map $c_{S,Q}:S\otimes Q\to Q\otimes S$ is given by some scalar $\lambda$, that is $c_{S,Q}=\lambda\tau_{S,Q}$. The tensor product in $\mathcal{C}$ is compatible with the $\mathcal{D}$-bimodule structure of $\mathcal{M}$, which in turn is compatible with the $\mathcal{D}$-bimodule structure of Vec, thus for $a$ the associativity constraint in $\mathcal{C}$ we have $a_{S,S,Q}=\text{id}_{S\otimes S\otimes Q}$, $a_{S,Q,S}=\text{id}_{S\otimes Q\otimes S}$ and $a_{Q,S,S}=\text{id}_{Q\otimes S\otimes S}$.
@@
$S\otimes Q\otimes S$$S\otimes S\otimes Q$$Q\otimes S\otimes S$$\mathbbm{1}\otimes Q$$Q\otimes\mathbbm{1}$$\scriptstyle{c_{S,Q}\otimes\text{id}_{S}}$$\scriptstyle{\text{id}_{S}\otimes c_{S,Q}}$$\scriptstyle{c_{S\otimes S,Q}}$$\scriptstyle{c_{\mathbbm{1},Q}}$
the upper triangle is commutative by the axiom of the braiding (suppressing the associativity constraints which are identities as explained above), and the lower square commutes by the naturality of the braiding. This diagram implies that $\lambda^{2}=1$, thus in particular $\lambda^{-1}=\lambda$.
Now consider the following diagram:
$S\otimes Q\otimes S$$Q\otimes S\otimes S$$Q\otimes S\otimes S$$S\otimes Q$$Q\otimes S$$\scriptstyle{c_{S,Q}\otimes\text{id}_{S}}$$\scriptstyle{\text{id}_{S}\otimes\psi}$$\scriptstyle{\text{id}_{Q}\otimes c_{S,S}}$$\scriptstyle{\psi\otimes\text{id}_{S}}$$\scriptstyle{c_{S,Q}}$
this diagram commutes due to the naturality of the braiding, so we get an equality:
$\lambda(\psi\otimes\text{id}_{S})\circ(\text{id}_{Q}\circ\tau_{S,S})\circ(\tau_{S,Q}\otimes\text{id}_{S})=(\lambda\tau_{S,Q})\circ(\text{id}_{S}\otimes\psi)$
which must imply that $\lambda=-\lambda$, giving a contradiction, thus $r=0$.
@@ 4.3. De-equivariantization of near-group categories
By 2.3.1 $\mathcal{C}_{H}$ is a well-defined braided finite tensor category, and the free module functor $F:\mathcal{C}\to\mathcal{C}_{H}$ is a braided tensor functor. This functor is surjective and normal by [BN3].
Let $Y\in\mathcal{C}_{H}$ be a simple object. By Lemma 2.1 in [BN2], $Y$ is a quotient of $F(X)$ for some $X\in\mathcal{C}$, that is we have a surjective morphism $\psi:F(X)\to Y$. Since $Y$ is simple and $F$ is a tensor functor (hence additive), we may assume that $X\in\mathcal{C}$ is indecomposable, and we can choose $X$ such that $\psi:F(X)\to L$ is non-zero and $X$ has minimal Loewy length. Let $L$ be the top of $X$, then $L=Q$ or $L\in\mathcal{D}$, and we have a surjective morphism $\varphi:X\to L$.
@@
If $L=Q$, by the assumption that $X$ is indecomposable, we must have $X=Q$. By [EO], surjective tensor functors map projective objects to projective objects, thus $F(X)=Q\otimes A$ is projective in $\textrm{Mod}_{\mathcal{C}}(A)$ for $A:=\text{Fun}(H,\mathbbm{k})$ the algebra object in $\text{Rep}(H)$. By Lemma 7.8.12 in [EGNO], $\text{Hom}_{A}(Q\otimes A,Y)=\text{Hom}_{\mathcal{C}}(Q,\bar{Y})$ for $\bar{Y}$ the $A$-module $Y$ considered as an object in $\mathcal{C}$, thus as an object in $\mathcal{C}$, we have $Y=Z\oplus Q^{\oplus m}$ for some $m\in\mathbb{N}$, $Z\in\mathcal{D}$.
@@
The algebra $A$ is separable (as an algebra object in $\text{Rep}(H)$), so $Y$ is a direct summand of $Y\otimes A$ in $\textrm{Mod}_{\mathcal{C}}(A)$, so to prove that $Y$ is projective in $\textrm{Mod}_{\mathcal{C}}(A)$, it is enough to prove that $Y\otimes A$ is projective in $\textrm{Mod}_{\mathcal{C}}(A)$, but this is clear because of Lemma 7.8.12 in [EGNO] and the fact that $\bar{Y}$ is a direct sum of copies of $Q$. In addition, it is clear from the proof that such $Y$ is uniquely determined by the fact that $L=Q$.
@@
Let $\mathcal{C}$ be a braided non-semisimple near-group category, i.e. a braided simple extension of a braided pointed finite tensor category $\mathcal{D}$. By 1, $\mathcal{C}$ is a $\text{Pic}(\mathcal{D})$-braided near-group category. By 3.2.2 and 3.2.6, the category $\mathcal{E}_{+}$ generated by objects in $\mathcal{D}$ that cetralize the unique simple projective object $Q\in\mathcal{C}$ is a symmetric subcategory of $\mathcal{D}$ which is central in $\mathcal{C}$. Hence, by 4.3.3, taking $H:=\text{Pic}(\mathcal{E}_{+})$ we get that $\mathcal{C}_{H}$ is a non-degenerate $G_{Q}$-braided near-group category and we get a modularization exact sequence:
@@ 4.4. Non-degenerate near-group categories
Let $L_{g}\in\mathcal{D}$ be a non-trivial simple object, and denote $s_{Q,L_{g}}=\lambda_{g}$, then . By Eq. 4 for the triplet $(Q,Q,L_{g})$ we get that $s_{Q\otimes Q,L_{g}}=\lambda_{g}^{2}\text{id}_{Q\otimes Q\otimes L_{g}}$. On the other hand, by Eq. 6, the naturality of the braiding, and the fact that $s_{Q\otimes Q,L_{g}}$ is a scalar map, we get that $s_{\bar{P},L_{g}}=\lambda_{g}^{2}\bar{P}\otimes L_{g}$. In particular, for $P_{e}$ the projective cover of $\mathbbm{1}$ we have $s_{P_{e},L_{g}}=\lambda_{g}^{2}\text{id}_{P_{e}\otimes L_{g}}$. $\mathcal{C}$ is non-degenerate, thus unimodular (by [EGNO]), so $\mathbbm{1}$ is the socle of $P_{e}$, thus $\lambda^{2}_{g}\text{id}_{L_{g}}=s_{\mathbbm{1},L_{g}}=\text{id}_{L_{g}}$, hence $\lambda_{g}^{2}=1$.
@@
Let $\mathcal{C}$ be a non-degenerate, braided non-semisimple near-group category with fusion rule $(G,0)$. By 4.4.1, we have $G=\mathbb{Z}_{2}$. This implies that $\mathcal{D}$ is a braided finite tensor category with only two simple objects, both invertible. Since $\mathcal{D}_{f}=\text{sVec}$ is a symmetric subcategory of $\mathcal{D}$, there is some Lagrangian subcategory $\mathcal{D}_{0}$ in $\mathcal{D}$, and we must have $\text{sVec}\subset\mathcal{D}_{0}$, so by [D], $\mathcal{D}_{0}$ is equivalent to a category of representations of some supergroup, and since $|G|=2$ this supergroup must be $\wedge W$ for some purely odd supervector space $W$, that is $\mathcal{D}_{0}\cong\textrm{sRep}(W)$. Now using [GS] we get the required result. ∎
@@
for $\bar{\mathcal{C}}$ a non-degenerate braided non-semisimple near-group category. By 3, $\bar{\mathcal{C}}$ is a simple extension of $\bar{\mathcal{D}}$ which is braid equivalent to $\textrm{sRep}(W\oplus W^{*})$. A direct calculation shows that $\text{\rm FPdim}(\bar{\mathcal{D}})=2^{2\dim(W)+1}$, so denoting by $Q_{0}$ the unique simple projective object of $\bar{\mathcal{C}}$, we have $\text{\rm FPdim}(Q_{0})=\sqrt{2}\cdot 2^{\dim(W)}$. Denoting by $Q$ the unique simple object in $\mathcal{C}$, we know that the functor $\mathcal{C}\to\bar{\mathcal{C}}$ is the free module functor and it maps $Q$ to $Q_{0}$. Since this is a surjective functor, by [EO] it preserve Frobenius-Perron dimensions, thus $\text{\rm FPdim}(Q)=\sqrt{2}\cdot^{\dim(W)}\notin\mathbb{N}$. ∎
@@
By 3, the category $\mathcal{C}_{H}$ has a braided subcategory $\tilde{\mathcal{D}}$ which is tensor equivalent to $\textrm{sRep}(W\oplus W^{*})$ for some purely odd supervector space $W$, and by [GS], $\tilde{S}\in\tilde{\mathcal{D}}^{\prime}$.

### v2

@@ 4.2. Non-semisimple braided near-group categories
Theorem 1. Let $\mathcal{C}$ be a non-semisimple braided near-group category with generalized fusion rule $(G,r)$, then $r=0$, and consequently $\mathcal{C}$ is weakly integral.
@@
Let $\mathcal{C}$ be a braided non-semisimple near-group category with a generalized fusion rule $(G,r)$, and assume that $r>0$. By 4.2.2, we get that $\mathcal{D}$ is a symmetric subcategory of $\mathcal{C}$ and $Q\in\mathcal{D}^{\prime}$. By [15], since $\mathcal{D}$ is not semisimple, it is braid equivalent to $\textrm{sRep}(G\ltimes W,u)$, a category of finite supermodules of the supergroup $G\ltimes W$, with $u\in G$ a central element of order $2$ acting as the parity. Since $\mathcal{D}$ is pointed, the simple objects of $\mathcal{D}$ generate a fusion subcategory, which is braid equivalent to $\text{Vec}_{G\times\mathbb{Z}/2\mathbb{Z}}$ (the central element $u$ generates $\mathbb{Z}/2\mathbb{Z}$). In particular, there exist a simple object $S\in\mathcal{D}_{f}\subset\mathcal{D}$ such that $S\otimes S=\mathbbm{1}$ and $c_{S,S}=-\tau_{S,S}$ for $\tau_{X,Y}$ the flip morphism $\tau_{X,Y}:X\otimes Y\to Y\otimes X$.
The tensor product in $\mathcal{C}$ is compatible with the left and right $\mathcal{D}$-module structure of $\mathcal{M}$, which in turn is compatible with the left and right $\mathcal{D}$-module structure of Vec, thus for $a$ the associativity constraint in $\mathcal{C}$ we have $a_{S,S,Q}=\text{id}_{S\otimes S\otimes Q}$ and $a_{Q,S,S}=\text{id}_{Q\otimes S\otimes S}$. The objects $S$ and $Q$ are simple, and $S$ is invertible, thus by abuse of notation, we can treat the morphisms $c_{S,S},c_{S,Q},c_{Q,S}$ and $a_{S,Q,S}$ as scalars.
@@
$S\otimes(Q\otimes S)$$(S\otimes Q)\otimes S$$S\otimes S\otimes Q$$Q\otimes S\otimes S$$\mathbbm{1}\otimes Q$$Q\otimes\mathbbm{1}$$\scriptstyle{a^{-1}_{S,Q,S}}$$\scriptstyle{c_{S,Q}\otimes\text{id}_{S}}$$\scriptstyle{\text{id}_{S}\otimes c_{S,Q}}$$\scriptstyle{c_{S\otimes S,Q}}$$\scriptstyle{c_{\mathbbm{1},Q}}$
the upper rectangle commutes by the axiom of the braiding (suppressing the associativity constraints $a_{Q,S,S}$ and $a_{S,S,Q}$ which are identities as explained above), and the lower square commutes by the naturality of the braiding. This diagram implies that $a_{S,Q,S}\cdot(c_{S,Q}^{2})=1$.
Now consider the following diagram. For ease of read, we use the notation $S^{\prime}$ to distinguish the second copy of $S$ from the first:
$S\otimes(S^{\prime}\otimes Q)$$(S\otimes S^{\prime})\otimes Q$$S\otimes(Q\otimes S^{\prime})$$(S^{\prime}\otimes S)\otimes Q$$(S\otimes Q)\otimes S^{\prime}$$S^{\prime}\otimes(S\otimes Q)$$(Q\otimes S)\otimes S^{\prime}$$(S^{\prime}\otimes Q)\otimes S$$Q\otimes(S\otimes S^{\prime})$$Q\otimes(S^{\prime}\otimes S)$$(Q\otimes S^{\prime})\otimes S$$\scriptstyle{a^{-1}_{S,S^{\prime},Q}}$$\scriptstyle{c_{S,S^{\prime}}\otimes\text{id}_{Q}}$$\scriptstyle{\text{id}_{S}\otimes c_{Q,S^{\prime}}}$$\scriptstyle{a_{S^{\prime},S,Q}}$$\scriptstyle{c_{S\otimes Q,S^{\prime}}}$$\scriptstyle{c_{S,Q}\otimes\text{id}_{S}^{\prime}}$$\scriptstyle{a_{S,Q,S^{\prime}}}$$\scriptstyle{a_{Q,S,S^{\prime}}}$$\scriptstyle{\text{id}_{S}^{\prime}\otimes c^{-1}_{Q,S}\circ a_{S^{\prime},Q,S}}$$\scriptstyle{\text{id}_{Q}\otimes c_{S,S^{\prime}}}$$\scriptstyle{a^{-1}_{Q,S^{\prime},S}}$$\scriptstyle{c_{Q,S^{\prime}}\otimes\text{id}_{S}}$
both rectangles commute by Eq. 1 for the triplet $(S,Q,S)$. From the upper rectangle, by Remark 3.2.1 for $S\otimes Q\cong Q$ and $c_{S,S^{\prime}}=-1$, we get that $a_{S,Q,S}=-1$, while the lower rectangle shows that $c_{S,Q}=c_{Q,S}$.
$a_{S,Q,S}=-1$ implies that $(c_{S,Q})^{2}=-1$, that is $c_{Q,S}^{2}=-1$, but $c_{Q,S}\cdot{c_{S,Q}}=1$ (because $S\in\mathcal{E}_{+}$), thus $c_{Q,S}=-c_{S,Q}$, which yields a contradiction.
@@ 4.3. De-equivariantization of near-group categories
By 2.3.1 $\mathcal{C}_{H}$ is a well-defined braided finite tensor category, and the free module functor $F:\mathcal{C}\to\mathcal{C}_{H}$ is a braided tensor functor. This functor is surjective and normal by [10].
Let $Y\in\mathcal{C}_{H}$ be a simple object. By Lemma 2.1 in [9], $Y$ is a quotient of $F(X)$ for some $X\in\mathcal{C}$, that is we have a surjective morphism $\psi:F(X)\to Y$. Since $Y$ is simple and $F$ is a tensor functor (hence additive), we may assume that $X\in\mathcal{C}$ is indecomposable, and we can choose $X$ such that $\psi:F(X)\to Y$ is non-zero and $X$ has minimal Loewy length. Let $L$ be the top of $X$, then $L=Q$ or $L\in\mathcal{D}$, and we have a surjective morphism $\varphi:X\to L$.
@@
If $L=Q$, by the assumption that $X$ is indecomposable, we must have $X=Q$. By [13], surjective tensor functors map projective objects to projective objects, thus $F(X)=Q\otimes A$ is projective in $\textrm{Mod}_{\mathcal{C}}(A)$ for $A:=\text{Fun}(H,\mathbbm{k})$ the algebra object in $\text{Rep}(H)$. By Lemma 7.8.12 in [1], $\text{Hom}_{A}(Q\otimes A,Y)=\text{Hom}_{\mathcal{C}}(Q,\bar{Y})$ for $\bar{Y}$ the $A$-module $Y$ considered as an object in $\mathcal{C}$, thus as an object in $\mathcal{C}$, we have $Y=Z\oplus Q^{\oplus m}$ for some $m\in\mathbb{N}$, $Z\in\mathcal{D}$.
@@
The algebra $A$ is separable (as an algebra object in $\text{Rep}(H)$), so $Y$ is a direct summand of $Y\otimes A$ in $\textrm{Mod}_{\mathcal{C}}(A)$, so to prove that $Y$ is projective in $\textrm{Mod}_{\mathcal{C}}(A)$, it is enough to prove that $Y\otimes A$ is projective in $\textrm{Mod}_{\mathcal{C}}(A)$, but this is clear because of Lemma 7.8.12 in [1] and the fact that $\bar{Y}$ is a direct sum of copies of $Q$. In addition, it is clear from the proof that such $Y$ is uniquely determined by the fact that $L=Q$.
@@
Let $\mathcal{C}$ be a braided non-semisimple near-group category, i.e. a braided simple extension of a braided pointed finite tensor category $\mathcal{D}$. By 1, $\mathcal{C}$ is a $\text{Pic}(\mathcal{D})$-braided near-group category. By 3.2.2 and 3.2.6, the category $\mathcal{E}_{+}$ generated by objects in $\mathcal{D}$ that centralize the unique simple projective object $Q\in\mathcal{C}$ is a symmetric subcategory of $\mathcal{D}$ which is central in $\mathcal{C}$. Hence, by 4.3.3, taking $H:=\text{Pic}(\mathcal{E}_{+})$ we get that $\mathcal{C}_{H}$ is a non-degenerate $G_{Q}$-braided near-group category and we get a modularization exact sequence:
@@ 4.4. Non-degenerate near-group categories
Let $L_{g}\in\mathcal{D}$ be a non-trivial simple object, and denote $s_{Q,L_{g}}=\lambda_{g}$. By Eq. 4 for the triplet $(Q,Q,L_{g})$ we get that $s_{Q\otimes Q,L_{g}}=\lambda_{g}^{2}\text{id}_{Q\otimes Q\otimes L_{g}}$. On the other hand, by Eq. 6, the naturality of the braiding, and the fact that $s_{Q\otimes Q,L_{g}}$ is a scalar map, we get that $s_{\bar{P},L_{g}}=\lambda_{g}^{2}\bar{P}\otimes L_{g}$. In particular, for $P_{e}$ the projective cover of $\mathbbm{1}$ we have $s_{P_{e},L_{g}}=\lambda_{g}^{2}\text{id}_{P_{e}\otimes L_{g}}$. $\mathcal{C}$ is non-degenerate, thus unimodular (by [1]), so $\mathbbm{1}$ is the socle of $P_{e}$, thus $\lambda^{2}_{g}\text{id}_{L_{g}}=s_{\mathbbm{1},L_{g}}=\text{id}_{L_{g}}$, hence $\lambda_{g}^{2}=1$.
@@
Let $\mathcal{C}$ be a non-degenerate, braided non-semisimple near-group category with fusion rule $(G,0)$. By 4.4.1, we have $G=\mathbb{Z}_{2}$. This implies that $\mathcal{D}$ is a braided finite tensor category with only two simple objects, both invertible. Since $\mathcal{D}_{f}=\text{sVec}$ is a symmetric subcategory of $\mathcal{D}$, there is some Lagrangian subcategory $\mathcal{D}_{0}$ in $\mathcal{D}$, and we must have $\text{sVec}\subset\mathcal{D}_{0}$, so by [15], $\mathcal{D}_{0}$ is equivalent to a category of representations of some supergroup, and since $|G|=2$ this supergroup must be $\wedge W$ for some purely odd supervector space $W$, that is $\mathcal{D}_{0}\cong\textrm{sRep}(W)$. Now using [8] we get the required result. ∎
@@
for $\bar{\mathcal{C}}$ a non-degenerate braided non-semisimple near-group category. By 3, $\bar{\mathcal{C}}$ is a simple extension of $\bar{\mathcal{D}}$ which is braid equivalent to $\textrm{sRep}(W\oplus W^{*})$. A direct calculation shows that $\text{\rm FPdim}(\bar{\mathcal{D}})=2^{2\dim(W)+1}$, so denoting by $Q_{0}$ the unique simple projective object of $\bar{\mathcal{C}}$, we have $\text{\rm FPdim}(Q_{0})=\sqrt{2}\cdot 2^{\dim(W)}$. Denoting by $Q$ the unique simple object in $\mathcal{C}$, we know that the functor $\mathcal{C}\to\bar{\mathcal{C}}$ is the free module functor and it maps $Q$ to $Q_{0}$. Since this is a surjective functor, by [13] it preserve Frobenius-Perron dimensions, thus $\text{\rm FPdim}(Q)=\sqrt{2}\cdot^{\dim(W)}\notin\mathbb{N}$. ∎
@@
By 3, the category $\mathcal{C}_{H}$ has a braided subcategory $\tilde{\mathcal{D}}$ which is tensor equivalent to $\textrm{sRep}(W\oplus W^{*})$ for some purely odd supervector space $W$, and by [8], $\tilde{S}\in\tilde{\mathcal{D}}^{\prime}$.